# RTNinja: a generalized machine learning framework for analyzing random telegraph noise signals in nanoelectronic devices

### v2

@@ V Scope of applicability
This section outlines the important conditions and constraints under which the RTNinja framework can be effectively utilized. While the framework is designed to be flexible and robust across a range of tasks, its performance and reliability are contingent upon specific assumptions and data characteristics. Understanding these constraints is critical for practitioners to determine whether the framework is appropriate for a given problem, and to anticipate potential limitations or challenges. The main guidelines that define the scope of applicability are as follows:
(i) Stationary signal: The framework assumes that the RTN signal remains stationary over the observation period. If device conditions drift over time or external interferences are present, the stationarity assumption may be violated, potentially compromising the accuracy of the analysis. To ensure reliable extraction of RTN parameters, such non-stationary components should be removed prior to applying the framework. A dedicated algorithm for drift correction has been recently reported and can be used in conjunction with this framework to enhance robustness [35].
@@
The RTNinja framework has already demonstrated its versatility and reliability through successful application to a variety of experimental RTN datasets across different device technologies. It has been effectively utilized for the analysis of nanoscale transistors and emerging device architectures, providing consistent and physically meaningful extraction of RTN parameters under diverse measurement conditions [36, 37, 38, 39]. These successful applications highlight the framework’s robustness and adaptability, while also validating the practical relevance of the assumptions outlined above. Collectively, they illustrate that within its defined scope, RTNinja serves as a powerful and generalizable tool for RTN signal analysis in advanced device technologies.
@@ VI Conclusion
In this work, we presented RTNinja, a generalized and fully automated machine learning framework for the analysis of random telegraph noise signals in nanoelectronic devices. By designing a modular architecture comprising the LevelsExtractor and SourcesMapper components, RTNinja effectively deconvolves complex RTN signals into their constituent sources and activity profiles without requiring prior assumptions about the number or nature of these sources. A key strength of RTNinja lies in its agnosticism to the underlying statistical distributions of source attributes and its independence from device-specific assumptions. This makes it broadly applicable across a wide range of nanoelectronic technologies and even beyond. The framework’s ability to operate without prior knowledge of the number or nature of RTN sources positions it as a powerful tool for both exploratory defect characterization and large-scale reliability benchmarking. Extensive Monte Carlo validation across 7000 datasets shows that RTNinja delivers accurate, scalable, and automated analysis. It reliably reconstructs RTN signals and extracts source parameters, even in challenging scenarios with overlapping levels, undersampled sources, or high noise floors. RTNinja unlocks new possibilities for applying RTN analysis in high-impact technology domains. In the semiconductor industry, it enables benchmarking of state-of-the-art devices — for instance, in dynamic random access memory components — by quantifying their susceptibility to defect-induced noise. This, in turn, guides circuit-level reliability assessments and memory architecture design. In quantum computing, where charge noise severely limits qubit coherence and fidelity, RTNinja can uniquely characterize individual fluctuations, offering critical insights for designing noise-resilient gate operations and control protocols. These capabilities make RTNinja a unique tool in RTN analysis, but also a versatile platform for shaping the development, qualification, and adoption of next-generation semiconductor and quantum technologies.
@@ Acknowledgements.
This work was part of the imec Industrial Affiliation Program. The authors thank Alexander Grill for helpful discussions during the RTNinja framework development.
@@
The synthetic datasets, RTNinja output files, the code for generating datasets, and the code for generating statistics from output files are available at 10.5281/zenodo.16750728. The code for RTNinja framework is currently part of a pending patent application (US 18/936,911; EP 23207819.6; CN 202411558787.3) and is therefore subject to legal restrictions.

### v3

@@ V Scope of applicability
This section outlines the important conditions and constraints under which the RTNinja framework can be effectively utilized. While the framework is designed to be flexible and robust across a range of tasks, its performance and reliability are contingent upon specific assumptions and data characteristics. Understanding these constraints is critical for practitioners to determine whether the framework is appropriate for a given problem and to anticipate potential limitations or challenges. The main guidelines that define the scope of applicability are as follows:
(i) Stationary signal: The framework assumes that the RTN signal remains stationary over the observation period. If device conditions drift over time or external interferences are present, the stationarity assumption may be violated, potentially compromising the accuracy of the analysis. To ensure reliable extraction of RTN parameters, such non-stationary components should be removed prior to applying the framework. A dedicated algorithm for drift correction has been recently reported and can be used in conjunction with this framework to enhance robustness [baseline].
@@
The RTNinja framework has already demonstrated its versatility and reliability through successful application to a variety of experimental RTN datasets across different device technologies. It has been effectively utilized for the analysis of nanoscale transistors and emerging device architectures, providing consistent and physically meaningful extraction of RTN parameters under diverse measurement conditions [piml, 3dnand, cryocmos, wearout]. These successful applications highlight the framework’s robustness and adaptability while also validating the practical relevance of the assumptions outlined above. Collectively, they illustrate that within its defined scope, RTNinja serves as a powerful and generalizable tool for RTN signal analysis in advanced device technologies.
@@ VI Conclusion
In this work, we presented RTNinja, a generalized and fully automated machine learning framework for the analysis of random telegraph noise signals in nanoelectronic devices. By designing a modular architecture comprising the LevelsExtractor and SourcesMapper components, RTNinja effectively deconvolves complex RTN signals into their constituent sources and activity profiles without requiring prior assumptions about the number or nature of these sources. A key strength of RTNinja lies in its agnosticism to the underlying statistical distributions of source attributes and its independence from device-specific assumptions. This makes it broadly applicable across a wide range of nanoelectronic technologies and even beyond. The framework’s ability to operate without prior knowledge of the number or nature of RTN sources positions it as a powerful tool for both exploratory defect characterization and large-scale reliability benchmarking. Extensive Monte Carlo validation across 7000 datasets shows that RTNinja delivers accurate, scalable, and automated analysis. It reliably reconstructs RTN signals and extracts source parameters, even in challenging scenarios with overlapping levels, undersampled sources, or high noise floors. RTNinja unlocks new possibilities for applying RTN analysis in high-impact technology domains. In the semiconductor industry, it enables benchmarking of state-of-the-art devices — for instance, in dynamic random access memory components — by quantifying their susceptibility to defect-induced noise. This, in turn, guides circuit-level reliability assessments and memory architecture design. In quantum computing, where charge noise severely limits qubit coherence and fidelity, RTNinja can uniquely characterize individual fluctuations, offering critical insights for designing noise-resilient gate operations and control protocols. These capabilities make RTNinja a unique tool in RTN analysis but also a versatile platform for shaping the development, qualification, and adoption of next-generation semiconductor and quantum technologies.
@@ Acknowledgements.
This work was part of the imec Industrial Affiliation Program. The authors thank Dr. Alexander Grill for helpful discussions during the RTNinja framework development.
@@
The synthetic datasets, RTNinja output files, the code for generating datasets, and the code for generating statistics from output files are available at 10.5281/zenodo.16750728. The code for the RTNinja framework is currently part of a pending patent application (US 18/936,911; EP 23207819.6; CN 202411558787.3) and is therefore subject to legal restrictions.